# Adaptive Error Correction for Entanglement Distillation

### v1

@@ 1. Step 1:
    Construct the $F_{\rm in}$ Sequence:
    In this step, we uniformly select 1000 points from $F_{\rm in}=0$ to $F_{\rm out}=1$. At the same time, we generate the probabilities $P_{X}$, $P_{Y}$, and $P_{Z}$ based on the characteristics of depolarizing noise.
@@ 2. Step 2:
    Construct the All Error List:
    We list all possible error events (including the case where no error occurs) and sort them in ascending order based on their error weight. That is, the first error in the list is the most probable error, the second error is the second most probable one, and so on. If we describe the error-correcting code using $\llbracket n,k,d\rrbracket$, then the total number of possible error combinations based on the syndrome is $2^{(n-k)}$.
@@ 3. Step 3:
    Write Data into the All-Error List:
    In this step, we need to perform three tasks. Using the points obtained from Step 1, calculate the probability of each error combination (for each of the 1000 selected points) and write this into the all-error list. Then, construct the Lookup Table and obtain the logical error based on the logical operator, then write it into the all-error list.
@@ 4. Step 4:
    Construct the Lookup Table:
    This step is performed simultaneously with Step 3. For each processed error combination, we obtain its corresponding syndrome. If this syndrome has not been recorded in the Lookup Table, we record the error combination that triggers this syndrome. If the syndrome has already been recorded, it means that the current error combination is not the smallest-weight error and does not need to be recorded, because in Step 2, we have already sorted the errors by weight.
@@ 5. Step 5:
    Obtain the Logical Error Corresponding to Each Error Combination:
    This step is performed simultaneously with Step 3. After Step 4, we use the logical operator to determine the logical error caused by the current error combination. Then, we retrieve the smallest weight error from the Lookup Table corresponding to this syndrome and test whether it can accurately predict the logical error caused by the current error combination: If it can predict correctly, record it as correctable by the Lookup Table. Otherwise, record it as uncorrectable. Finally, write the results into the error list.
@@ 6. Step 6:
    Generate the $F_{\rm in}$-$F_{\rm out}$ Map:
    From the all-error list, read all correctable logical error probabilities and accumulate them for each of the 1000 points (generated in Step 1). In the end, we obtain the $F_{\rm in}$-$F_{\rm out}$ map for 1000 points spanning $F_{\rm in}=0$ to $F_{\rm in}=1$.
@@ 1. Step 1:
    Construct the $F_{\rm in}$ sequence:
    In this step, we uniformly select a required number of points from $F_{\rm in}=0$ to $F_{\rm in}=1$ as needed.
@@ 2. Step 2:
    Perform error correction in this round:
    Based on the $F_{\rm in}$ points constructed in Step 1 and the chosen $\llbracket n,k,d\rrbracket$ error-correcting code, we use interp1() to look up the corresponding $F_{\rm in}$-$F_{\rm out}$ map in 0-repeater case, where the $F_{\rm out}$ value represents the post-error-correction $F_{\rm out}$ at this round.
@@ 3. Step 3:
    Compute the output fidelity of entanglement swap:
    Based on our discussion in the background (Eqn. (12)), since in this case each round of error correction is based on the same codes, with the same input fidelity at the beginning, we can simplify the formula to the following form:
    $F_{\rm eff}=\frac{1}{4}+\frac{3}{4}\left(\frac{4F_{\rm in}-1}{3}\right)^{n_{\rm QS}+1}$, where $n_{\rm QS}$ represents the total number of quantum swaps required in this round. The result obtained from this formula becomes the $F_{\rm in}$ for the next round.
@@ 4. Step 4:
    Obtain the final $F_{\rm in}$-$F_{\rm out}$ map:
    By repeating Step 2 and Step 3 with different codes in each round, we complete error correction across all rounds and obtain the final $F_{\rm in}$-$F_{\rm out}$ map.
Figure 7: the $F_{\rm in}$-$F_{\rm out}$ map for the 1-repeater case. This figure only shows the relationship between input fidelity and output fidelity based on different protocols. No coding means that no error-correcting code is used. For example, when $n=1$, $F_{\rm out}^{\text{No Coding}}=F_{\rm in}$,which means that when the input is a single qubit, the output fidelity corresponds to the $n=1$ qubit case. In general from $F_{\rm out}^{\text{No Coding}}=F_{\rm in}^{n}$.
@@ IV-A Goal
In Fig. 7, we observe that the choice of coding protocol can significantly affect the fidelity. We also observed that the logical qubit rate and distillable entanglement can be affected by different protocols. This implies that it is impossible to achieve an overall optimal result using a single code. Hence, our goal is to develop an adaptive strategy to switch between protocols at relatively optimal thresholds, balancing the trade-off between rate, fidelity, and distillable entanglement.
IV-B Results
@@
IV-B1 Single Repeater
@@
These switching points can be clearly observed in Fig. 8. The corresponding results for $R_{\rm out},F_{\rm out}$ and $D_{\rm out}$ are shown in Figs. 9.a10.a and 11.a, respectively.
9.a The $F_{\rm in}$-$R_{out}$ map for the 1-repeater case with optimal envelope.
10.a The $F_{\rm in}$-$F_{\rm out}$ map for the 1-repeater case with optimal envelope.
11.a The $F_{\rm in}$-$D_{out}$ map for the 1-repeater case with optimal envelope.

### v2

@@ 1. Step 1:
    Construct the $F_{\rm in}$ Sequence: In this step, we uniformly select 1000 points from $F_{\rm in}=0$ to $F_{\rm in}=1$. At the same time, we generate the probabilities $P_{X}$, $P_{Y}$, and $P_{Z}$ based on the characteristics of depolarizing noise.
@@ 2. Step 2:
    Construct the All Error List:We list all possible error events (including the case where no error occurs) and sort them in ascending order based on their error weight. That is, the first error in the list is the most probable error, the second error is the second most probable one, and so on. If we describe the error-correcting code using $\llbracket n,k,d\rrbracket$, then the total number of possible error combinations based on the syndrome is $2^{(n-k)}$.
@@ 3. Step 3:
    Write Data into the All-Error List: In this step, we need to perform three tasks. Using the points obtained from Step 1, calculate the probability of each error combination (for each of the 1000 selected points) and write this into the all-error list. Then, construct the Lookup Table and obtain the logical error based on the logical operator, then write it into the all-error list.
@@ 4. Step 4:
    Construct the Lookup Table: This step is performed simultaneously with Step 3. For each processed error combination, we obtain its corresponding syndrome. If this syndrome has not been recorded in the Lookup Table, we record the error combination that triggers this syndrome. If the syndrome has already been recorded, it means that the current error combination is not the smallest-weight error and does not need to be recorded, because in Step 2, we have already sorted the errors by weight.
@@ 5. Step 5:
    Obtain the Logical Error Corresponding to Each Error Combination: This step is performed simultaneously with Step 3. After Step 4, we use the logical operator to determine the logical error caused by the current error combination. Then, we retrieve the smallest weight error from the Lookup Table corresponding to this syndrome and test whether it can accurately predict the logical error caused by the current error combination: If it can predict correctly, record it as correctable by the Lookup Table. Otherwise, record it as uncorrectable. Finally, write the results into the error list.
@@ 6. Step 6:
    Generate the $F_{\rm in}$-$F_{\rm out}$ Map: From the all-error list, read all correctable logical error probabilities and accumulate them for each of the 1000 points (generated in Step 1). In the end, we obtain the $F_{\rm in}$-$F_{\rm out}$ map for 1000 points spanning $F_{\rm in}=0$ to $F_{\rm in}=1$.
Figure 6: the $F_{\rm in}$-$F_{\rm out}$ map for the 1-repeater case. This figure only shows the relationship between input fidelity and output fidelity based on different protocols. No coding means that no error-correcting code is used. For example, when $n=1$, $F_{\rm out}^{\text{No Coding}}=F_{\rm in}$,which means that when the input is a single qubit, the output fidelity corresponds to the $n=1$ qubit case. In general, $F_{\rm out}^{\text{No Coding}}=F_{\rm in}^{n}$.
@@ 1. Step 1:
    Construct the $F_{\rm in}$ sequence: In this step, we uniformly select a required number of points from $F_{\rm in}=0$ to $F_{\rm in}=1$ as needed.
@@ 2. Step 2:
    Perform error correction in this round: Based on the $F_{\rm in}$ points constructed in Step 1 and the chosen $\llbracket n,k,d\rrbracket$ error-correcting code, we use interp1() to look up the corresponding $F_{\rm in}$-$F_{\rm out}$ map in 0-repeater case, where the $F_{\rm out}$ value represents the post-error-correction $F_{\rm out}$ at this round.
@@ 3. Step 3:
[…]
@@ 4. Step 4:
    Obtain the final $F_{\rm in}$-$F_{\rm out}$ map: By repeating Step 2 and Step 3 with different codes in each round, we complete error correction across all rounds and obtain the final $F_{\rm in}$-$F_{\rm out}$ map.
@@ IV-A Goal
In Fig. 6, we observe that the choice of coding protocol can significantly affect the fidelity. We also observe that the logical qubit rate and distillable entanglement can be affected by different protocols. This implies that it is impossible to achieve an overall optimal result using a single code. Hence, our goal is to develop an adaptive strategy to switch between protocols at relatively optimal thresholds, balancing the trade-off between rate, fidelity, and distillable entanglement.
IV-B Efficiency Function
@@
To calculate the protocol rate, we must ensure that the number of output Bell pairs from the previous round matches the input required by the code used in the next round. Therefore, the rate must be calculated iteratively on a round-by-round basis.
IV-B1 First Round Using the $\llbracket n_{1},k_{1},d_{1}\rrbracket$ Code
When discussing the first round, we need to consider the number of repeaters ($n_{R}$) to be used. The corresponding input and output qubits are shown below.
$N^{(1)}_{\rm in}$ $=(n_{R}+1)\cdot n_{1}$ (17)
$K^{(1)}_{\rm out}$ $=k_{1}$ (18)
IV-B2 Second Round Using the $\llbracket n_{2},k_{2},d_{2}\rrbracket$ Code
The key point in the second round is to match $K^{(1)}_{\rm out}$ with $n_{2}$. Therefore, we use $L^{(1)}=\mathrm{lcm}(K^{(1)}_{\rm out},n_{2})$.Thus, the corresponding input and output qubits are shown below.
$N^{(2)}_{\rm in}$ $=\frac{L^{(1)}}{K^{(1)}_{\rm out}}\cdot N^{(1)}_{\rm in}$ (19)
$K^{(2)}_{\rm out}$ $=\frac{L^{(1)}}{n_{2}}\cdot k_{2}$ (20)
IV-B3 Third Round Using the $\llbracket n_{3},k_{3},d_{3}\rrbracket$ Code
Similar to the second round, we need to use $L^{(2)}=\mathrm{lcm}(K^{(2)}_{\rm out},n_{3})$ to ensure proper matching between them. Thus, the final input and output qubits are shown below.
$N^{(3)}_{\rm in}$ $=\frac{L^{(2)}}{K^{(2)}_{\rm out}}\cdot N^{(2)}_{\rm in}$ (21)
$K^{(3)}_{\rm out}$ $=\frac{L^{(2)}}{n_{3}}\cdot k_{3}$ (22)
Here, $N^{(3)}_{\rm in}$ and $K^{(3)}_{\rm out}$ aare the total number of Bell pairs required for the input and the total number of output Bell pairs, respectively. Thus, we can set $n_{in}=N^{(3)}_{\rm in}$ and $n_{out}=K^{(3)}_{\rm out}$.
IV-C Results
IV-C1 Single Repeater
@@
These switching points can be clearly observed in Fig. 7. The corresponding results for $R_{\rm out},F_{\rm out}$ and $D_{\rm out}$ are shown in Figs. $8.a$ and 8.a, respectively.
8.a The $F_{\rm in}$-$R_{out}$ map for the 1-repeater case with optimal envelope.
8.a The $F_{\rm in}$-$F_{\rm out}$ map for the 1-repeater case with optimal envelope.
8.a The $F_{\rm in}$-$D_{out}$ map for the 1-repeater case with optimal envelope.